OpenStudy (anonymous):

Two cars A and B are 200km apart and they are traveling toward each other at the speeds of 28km/h and 22 km/h respectively. When and where will they meet? 2. If the car A and B in problem 1 start from the same point, going toward the same direction, at what time will they be 18km apart? 3. If B starts at 1:00pm and gives chase starting at 1:15 from the same point as B, when and where will A overtake B? Thank you so much for answering!!

OpenStudy (lgbasallote):

like i was saying before distance traveled by car A + distance traveled by car B = total distance. do you agree?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

so what will be the total distance in this case?

OpenStudy (anonymous):

d= 50kph?

OpenStudy (lgbasallote):

kph means kilometers per hour. this is a unit of measurement for speed

OpenStudy (anonymous):

oh.. right it should be 50km?

OpenStudy (lgbasallote):

is there a given 50 km in the problem?

OpenStudy (anonymous):

no

OpenStudy (lgbasallote):

which one is given in km?

OpenStudy (anonymous):

is it 200km

OpenStudy (lgbasallote):

yes that's the total distance so distance traveled by car A + distance traveled by car B = 200

OpenStudy (lgbasallote):

so now we need to find the distance traveled by car A and B it is given that car A is traveling at 28 km/h we know that distance = speed x time let t = time so distance traveled by car A is 28t

OpenStudy (lgbasallote):

can you now find the distance traveled by car B?

OpenStudy (anonymous):

ok

OpenStudy (lgbasallote):

hint their time is the same

OpenStudy (anonymous):

so the distance of Car B = 22t?

OpenStudy (lgbasallote):

yup

OpenStudy (lgbasallote):

so now add 28t + 22t = 200 solve for t

OpenStudy (anonymous):

\[28t +22t = 200\] \[50t \div50 = 200 \div 50\] \[t = 4\]

OpenStudy (lgbasallote):

right so time is 4 hours

OpenStudy (lgbasallote):

they'll meet after 4 hours

OpenStudy (lgbasallote):

as for where they'll meet i dont quite get what it wants

OpenStudy (anonymous):

why not? how should it be?

OpenStudy (lgbasallote):

i dont know. like i said i dont know what it wants

OpenStudy (lgbasallote):

an obvious answer is they'll meet at 0 km lol

OpenStudy (anonymous):

ok then, let's proceed

OpenStudy (lgbasallote):

part 2 is exactly the same

OpenStudy (anonymous):

ok, but is it done

OpenStudy (anonymous):

how?

OpenStudy (lgbasallote):

what's done?

OpenStudy (anonymous):

* ok, but how is it done how to do part 2?

OpenStudy (lgbasallote):

same way distance traveled by car A + distance traveledby car B = total distance

OpenStudy (anonymous):

\[28t + 22t = 18km\] ??? is that how i should start it?

OpenStudy (lgbasallote):

yup

OpenStudy (lgbasallote):

uhmm wait...

OpenStudy (anonymous):

how do i get the Distance of A? okkk

OpenStudy (lgbasallote):

sorry should be subtraction..because its difference in speed

OpenStudy (lgbasallote):

28t - 22t = 18

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so t= 3hrs?

OpenStudy (lgbasallote):

yup

OpenStudy (anonymous):

then. what's next?

OpenStudy (lgbasallote):

for #3 it is unclear to me...can you rewrite it

OpenStudy (anonymous):

Right.. sorry bout that. If B starts at 1:00pm and A gives chase starting at 1:15 from the same point as B, when and where will A overtake B?

OpenStudy (lgbasallote):

so if car b starts at 1 pm then let his time be t car a starts at 1:15 so its time is t+15 agree?

OpenStudy (lgbasallote):

uhmm sorry should be time of car A should be t-15 because it starts 15 minutes later than car B

OpenStudy (anonymous):

ok

OpenStudy (lgbasallote):

so when car A overtakes car B then their distance traveled will be the same agree>

OpenStudy (anonymous):

yup

OpenStudy (lgbasallote):

so do you have an idea what the equation will look like?

OpenStudy (anonymous):

so should it be t= t - 15 ??

OpenStudy (lgbasallote):

nope..that would just cancel out the t

OpenStudy (lgbasallote):

remember distance traveled by car A = distance traveled by car B

OpenStudy (anonymous):

22t = 28t - 15 ??

OpenStudy (lgbasallote):

close

OpenStudy (lgbasallote):

remember t - 15 is the time itself.. should be 22t = 28(t-15)

OpenStudy (lgbasallote):

it makes a significant difference

OpenStudy (anonymous):

oh.. ok

OpenStudy (anonymous):

then we solve for t: ??

OpenStudy (lgbasallote):

yes

OpenStudy (anonymous):

22t = 28t - 420 420 = 28t - 22t 420 = 6t --- --- 6 6 t = 70hrs ???

OpenStudy (lgbasallote):

seems right

OpenStudy (anonymous):

so A will overtake be after 70 hrs ??

OpenStudy (anonymous):

where will A overtake B??/

OpenStudy (lgbasallote):

just substitute 28(70-15)

OpenStudy (anonymous):

ok so D = 1,540 km

OpenStudy (anonymous):

is D = 1,540 km ???

mathslover (mathslover):

Yes u r right ..@alynamarie D= 1540 km

OpenStudy (anonymous):

@mathslover thanks